If we want to guess the proportion of children who are blind to be within 2 percentage points of the sample proportion (ME=2%) with 95% confidence, how many children should we sample?
Answer #1

Solution :

Given that,

= 1 - = 0.5

margin of error = E = 0.02

Z/2 = 1.96

sample size = n = (Z / 2 / E )2 * * (1 - )

= (1.96 / 0.02)2 * 0.5 * 0.5

= 2401

sample size = 2401